Circular No. 7214

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cz IN NGC 5078 A. Williams, R. Martin, and S. Woodings report their discovery of an

apparent supernova (mag ∼ 16) on an image taken on June 21.53 UT in the course of the Perth Automated Supernova Search. A confirming image was obtained on June 23.4, while no object was present at this location brighter than mag ∼ 18 on May 22.6, and an earlier image shows no object to a limiting magnitude of ∼ 18.5. C. Livingston (Curtin University) and J. Biggs and A. Verveer (Perth Observatory) provide the following position for the supernova candidate from an unfiltered CCD image obtained around July 1.493 with the 0.25-m Mike Candy Telescope at Perth Observatory: α = 13^{h}19^{m}48^{s}.53 ± 0^{s}.02, δ = −27^{o}24′04′′.5 ± 0′′.2 (equinox 2000.0; mag 16.5), which is ∼ 36′′ northwest of the center of NGC 5078.

GRB 990704 A. Maury and B. Gaillard, Observatoire de la Côte d’Azur (OCA);

and M. Boer, Centre d’Etude Spatiale des Rayonnements, CESR/CNRS, report on behalf of the OCA/CESR GRB collaboration: “We observed the error box of GRB 990704 (α = 12^{h}19^{m}30^{s}, δ = −3^{o}48′.2, error-circle radius ∼ 7′; GCN 360, 366) with the OCA 0.90-m Schmidt telescope (+ unfiltered CCD). An object that is not present on the Palomar Sky Survey has been detected at α = 12^{h}19^{m}29^{s}.29, δ = −3^{o}47′25′′.8 at mag 19.4 on July 4.876 UT. Since the field was very low in the sky, the position reported here may be affected by somewhat large errors. Images of the field and of the optical transient are available at the following URL: http://wwwrc.obs-

azur.fr/schmidt/observations/GRB990704.html. For further informa-
tion contact either Michel.Boer@cesr.fr or maury@obs-azur.fr.

VARIABLE STAR TOWARDS UGC 11289 M. Modjaz and W. D. Li, University of California at Berkeley, on be-

half of the Lick Observatory Supernova Search, report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of a variable star in the field of UGC 11289. The possible cataclysmic variable, discovered and confirmed on unfiltered images taken on July 1.4 (mag ∼ 17.8) and 2.4 UT (mag ∼ 18.9), is located α = 18^{h}35^{m}56^{s}.27, δ = +22^{o}26′40′′.0 (equinox

2000.0; finding chart http://astron.berkeley.edu/~bait/1999/cv 1.

html). A KAIT image of the same field on June 27.4 (limiting mag 19.0) and 25.4 (limiting mag 19.5) showed nothing at this position.

1999 July 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green